# Category $\mathcal{O}$ for hybrid quantum groups and non-commutative Springer resolutions

### v3

@@ 4.2.1. Equivalence for Steinberg block
We can adjust the $\Lambda$-grading on $M_{\chi_{-\rho}}$ such that the isomorphism (4.7) respects the $\Lambda$-gradings on both sides. Therefore, the equivalence (4.6) is given by
@@
For any commutative Noetherian $S$-algebra $R$, the equivalence ${\mathfrak{G}}^{b}$ induces an equivalence
${\mathfrak{G}}^{b}:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{R}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{R},$
where ${\mathfrak{b}}_{R}:={\mathfrak{b}}_{S}\times_{\mathrm{Spec}S}\mathrm{Spec}R$.
Since the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}})$ is ${\mathbb{C}}[{\mathfrak{g}}]$-linear and $G$-equivariant, it induces a functor $\pi_{*}:{\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}\rightarrow\mathrm{Coh}^{T}({\mathfrak{b}}_{S})$. From Proposition 3.6 we deduce the following
@@
There is a commutative diagram
[…]
Recall that by [36, §3.1.2] there is a subalgebra ${\mathfrak{U}}^{\mathrm{hb}}_{\zeta}=Z_{\mathrm{Fr}}^{-}\otimes{\mathfrak{U}}^{0}_{\zeta}\otimes U^{+}_{\zeta}$ of $U^{\mathrm{hb}}_{\zeta}$, equipped with a Frobenius map
@@
compatible with the quantum Frobenius map $U^{+}_{\zeta}\rightarrow U{\mathfrak{n}}$, and the isomorphism $Z^{-}_{\mathrm{Fr}}\otimes{\mathfrak{U}}^{0}_{\zeta}\rightarrow{\mathbb{C}}[B\times_{T}T]$ by (2.1), where $T\rightarrow T$ is by the $l$-th power. The map $\mathrm{Fr}$ is a homomorphism of $\Lambda$-graded algebras, where we view ${\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}}$ as an $l\Lambda$-graded ring. In loc. cit. we defined a functor for any deformation ring $R$ of $U^{\mathrm{hb}}_{\zeta}$,
@@
$\mathrm{ind}_{c}:=({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{R}]}-:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{R})\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}.$
@@
Note that $\mathrm{ind}_{r}$ preserves the block decomposition.
@@
There is a commutative diagram
[…]
@@
4.2.2. Functors ${\mathbb{V}}^{b}_{r}$ and ${\mathbb{V}}^{b}_{c}$
Recall that ${\mathrm{A}}=\mathrm{End}^{\mathrm{op}}_{\tilde{{\mathfrak{g}}}}({\mathcal{E}})$ for some tilting bundle ${\mathcal{E}}$ on $\tilde{{\mathfrak{g}}}$. By [12, Lem 2.5.3], ${\mathcal{E}}$ has ${\mathcal{O}}_{\tilde{{\mathfrak{g}}}}$ as a direct summand. It follows that ${\mathrm{A}}$ is an algebra over $\Gamma({\mathcal{O}}_{\tilde{{\mathfrak{g}}}})={\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$, and that ${\mathrm{A}}$ has ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ as a direct summand as a ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-module. Moreover, the subalgebra ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ is central in ${\mathrm{A}}$. Hence ${\mathrm{A}}_{S}^{b}$ is an algebra over ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ and ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]\subset Z({\mathrm{A}}_{S}^{b})$. We have ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]^{T}={\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}$. Therefore, we have an algebra homomorphism
(4.8) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z({\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}).$
Note that the functor $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}})$ is represented by the $({\mathbb{C}}[{\mathfrak{g}}],{\mathrm{A}})$-bimodule $\Gamma(\tilde{{\mathfrak{g}}},{\mathcal{E}})$. Hence $\pi_{*}$ can be enhanced to $\pi_{*}:{\mathrm{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$ and induces a functor
${\mathbb{V}}^{b}_{c}:\ {\mathrm{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}\rightarrow\mathrm{Coh}^{T}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
by base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$.
@@ 4.2.2. Functors ${\mathbb{V}}^{b}_{r}$ and ${\mathbb{V}}^{b}_{c}$
Recall that by §2.3.3 there is an algebra homomorphism
@@
${\mathbb{V}}^{b}_{r}:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{T^{-\rho}_{0}}{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
By the identification (3.8) and the compatibility of ${\mathfrak{F}}^{\prime}$ and ${\mathfrak{F}}^{b}$, the two maps (4.8) and (4.9) coincide under the equivalence ${\mathfrak{F}}^{b}$. Therefore, we obtain the following enhancement of Proposition 4.5.
Corollary 4.7.
There is a commutative diagram
[…]

### v4

@@ 4.2. Equivalence ${\mathfrak{G}}^{b}$ for Steinberg block
We can adjust the $\Lambda$-grading on $M_{\chi_{-\rho}}$ such that the isomorphism (4.7) respects the $\Lambda$-gradings (in completion sense) on both sides. Therefore, the equivalence (4.6) is given by
@@
In particular, we have an isomorphism ${\mathfrak{G}}^{b}({\mathbb{C}}[{\mathfrak{b}}_{S}]\otimes{\mathbb{C}}_{\lambda})=M(-\rho+l\lambda)_{S}$ for any $\lambda\in\Lambda$.
By (2.5) the equivalence ${\mathfrak{G}}^{b}$ induces an equivalence
${\mathfrak{G}}^{b}:\ \mathrm{Coh}^{T}({\mathfrak{b}}_{R}){\ \xrightarrow{\sim}\ }{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{R}$
for any commutative Noetherian $S$-algebra $R$, where ${\mathfrak{b}}_{R}:={\mathfrak{b}}_{S}\times_{\mathrm{Spec}S}\mathrm{Spec}R$.
[…]
$\pi_{*}:{\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{K}\rightarrow\mathrm{Coh}^{K}({\mathfrak{b}}_{S})$
which is represented by the $({\mathbb{C}}[{\mathfrak{b}}_{S}],{\mathsf{A}}_{S}^{b})$-bimodule $\Gamma(\widetilde{{\mathfrak{g}}},{\mathcal{E}})\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}]$. The functor $\pi_{*}$ admits a right adjunction
$\pi^{!}:=\mathrm{Hom}_{{\mathbb{C}}[{\mathfrak{b}}_{S}]}(\Gamma(\widetilde{{\mathfrak{g}}},{\mathcal{E}})\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}],-):\ \mathrm{Coh}^{K}({\mathfrak{b}}_{S})\rightarrow{\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{K}.$
From Proposition 3.6 we deduce the following
@@
There are commutative diagrams
(4.8) ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$$\mathrm{Coh}^{T}({\mathfrak{b}}_{S})$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S},$$\scriptstyle{{\mathfrak{F}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{\pi_{*}}$$\scriptstyle{{\mathsf{T}}^{-\rho}_{0}}$$\scriptstyle{{\mathfrak{G}}^{b}}\scriptstyle{\simeq}$
and
(4.9) ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$$\mathrm{Coh}^{T}({\mathfrak{b}}_{S})$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S}.$$\scriptstyle{{\mathfrak{F}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{{\mathfrak{G}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{\pi^{!}}$$\scriptstyle{{\mathsf{T}}_{-\rho}^{0}}$
4.3. Induction functors
Recall that by [34, §3.1.2] there is a subalgebra ${\mathfrak{U}}^{\mathrm{hb}}_{\zeta}=Z_{\mathrm{Fr}}^{-}\otimes{\mathfrak{U}}^{0}_{\zeta}\otimes U^{+}_{\zeta}$ of $U^{\mathrm{hb}}_{\zeta}$, equipped with a Frobenius map
@@ 4.3. Induction functors
which is compatible with the quantum Frobenius map $U^{+}_{\zeta}\rightarrow U{\mathfrak{n}}$, and with the isomorphism $Z^{-}_{\mathrm{Fr}}\otimes{\mathfrak{U}}^{0}_{\zeta}{\ \xrightarrow{\sim}\ }{\mathbb{C}}[B\times_{T}T]$ induced by (2.1), where $T\rightarrow T$ is by the $l$-th power. The map $\mathrm{Fr}$ is a homomorphism of $\Lambda$-graded algebras, where we view ${\mathbb{C}}[B\times_{T}T]\rtimes U{\mathfrak{n}}$ as an $l\Lambda$-graded ring. In loc. cit. we defined a functor for any deformation ring $R$ of $U^{\mathrm{hb}}_{\zeta}$,
@@
$\mathrm{ind}_{c}:=({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\otimes_{{\mathbb{C}}[{\mathfrak{b}}_{R}]}-:\ \mathrm{QCoh}^{T}({\mathfrak{b}}_{R})\rightarrow({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}.$
@@
Note that $\mathrm{ind}_{r}$ preserves the block decomposition in §2.3.3.
@@
Let $R$ be a commutative Noetherian $S$-algebra. There is a commutative diagram
(4.10) $\mathrm{Coh}^{T}({\mathfrak{b}}_{R})$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{R}$$({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$$U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}.$$\scriptstyle{{\mathfrak{G}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{\mathrm{ind}_{c}}$$\scriptstyle{\mathrm{ind}_{r}}$$\scriptstyle{{\mathrm{I}}^{\mathrm{hb}}_{1}}$
@@
Let $R$ be a commutative Noetherian $S$-algebra. Let $\nu\in\Lambda$. We denote the truncation functor on $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{mod}^{\Lambda}_{R}$ by the poset ideal $\{\lambda\in\Lambda|\lambda\leq\nu\}$ (see §2.3.1) as
$\tau_{r}^{\leq\nu}:\ U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda}_{R}\rightarrow U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\leq\nu}_{R}.$
We denote by
${\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B,\leq\nu}$
the category of modules in $({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$ whose $T$-weights are $\leq\nu$, which is a subcategory in ${\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B}$ (since the $U{\mathfrak{n}}$-action on a module with $T$-weights $\leq\nu$ is locally unipotent). By similar argument as in §2.3.1, there is a truncation functor
$\tau^{\leq\nu}_{c}:\ ({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}\rightarrow{\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B,\leq\nu}$
by taking the maximal quotient whose $T$-weights are $\leq\nu$, which is left adjoint to the natural inclusion. We abbreviate ${\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{mod}^{B,\leq\nu}:={\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{Mod}^{B,\leq\nu}\cap{\mathbb{C}}[{\mathfrak{b}}_{R}]\text{-}\mathrm{mod}^{B}$.
Lemma 4.7.
There is a natural isomorphism of functors
(4.11) ${\mathrm{I}}^{\mathrm{hb}}_{1}\circ\tau^{\leq\nu}_{c}=\tau^{\leq-\rho+l\nu}_{r}\circ{\mathrm{I}}^{\mathrm{hb}}_{1}.$
from $({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}})\text{-}\mathrm{Mod}^{T}$ to $U^{\mathrm{hb}}_{\zeta}\text{-}\mathrm{Mod}^{\leq-\rho+l\nu}_{R}$.
Proof.
By the right exactness of the functors, it is enough to construct natural isomorphism on the projective generators ${\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}}\otimes{\mathbb{C}}_{\lambda}$, with $\lambda\in\Lambda$. We have an isomorphism of $\Lambda$-graded $U^{\mathrm{hb}}_{\zeta}\otimes R$-modules
${\mathrm{I}}^{\mathrm{hb}}_{1}({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}}\otimes{\mathbb{C}}_{\lambda})=U^{\mathrm{hb}}_{\zeta}\otimes_{U^{\mathrm{hb},\geq}_{\zeta}}(U{\mathfrak{n}}\otimes R_{-\rho+l\lambda}),$
where $U{\mathfrak{n}}\otimes R_{\mu}$ is viewed as a $U^{\mathrm{hb},\geq}_{\zeta}-$module via the map $U^{+}_{\zeta}\otimes{\mathfrak{U}}^{0}_{\zeta}\xrightarrow{\mathrm{Fr}\otimes\tau_{\mu}}U{\mathfrak{n}}\otimes{\mathfrak{U}}^{0}_{\zeta}\rightarrow U{\mathfrak{n}}\otimes R$, for any $\mu\in\Lambda$. Hence we have isomorphisms
${\mathrm{I}}^{\mathrm{hb}}_{1}\big(\tau^{\leq\nu}_{c}({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}}\otimes{\mathbb{C}}_{\lambda})\big)$ $={\mathrm{I}}^{\mathrm{hb}}_{1}\big({\mathbb{C}}[{\mathfrak{b}}_{R}]\otimes\big(U{\mathfrak{n}}/\bigoplus_{\eta\nleq\nu-\lambda}(U{\mathfrak{n}})_{\eta}\big)\otimes{\mathbb{C}}_{\lambda}\big)$
$=U^{\mathrm{hb}}_{\zeta}\otimes_{U^{\mathrm{hb},\geq}_{\zeta}}(U{\mathfrak{n}}/\bigoplus_{\eta\nleq\nu-\lambda}(U{\mathfrak{n}})_{\eta}\big)\otimes R_{-\rho+l\lambda})$
$=\tau^{\leq-\rho+l\nu}_{r}\big({\mathrm{I}}^{\mathrm{hb}}_{1}({\mathbb{C}}[{\mathfrak{b}}_{R}]\rtimes U{\mathfrak{n}}\otimes{\mathbb{C}}_{\lambda})\big).\qed$
4.4. Functors ${\mathbb{V}}^{b}$
@@
Recall that ${\mathsf{A}}=\mathrm{End}^{\mathrm{op}}_{\widetilde{{\mathfrak{g}}}}({\mathcal{E}})$ for some tilting bundle $\mathcal{E}$} on $\widetilde{{\mathfrak{g}}}$. By [11, Lem 2.5.3], $\mathcal{E}$} has ${\mathcal{O}}_{\widetilde{{\mathfrak{g}}}}$ as a direct summand. It follows that $\mathsf{A}$} is an algebra over $\Gamma({\mathcal{O}}_{\widetilde{{\mathfrak{g}}}})={\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$, and that $\mathsf{A}$} has ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ as a direct summand as a ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$-module. Moreover, the subalgebra ${\mathbb{C}}[{\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$ is central in $\mathsf{A}$}. Hence ${\mathsf{A}}_{S}^{b}$ contains a central subalgebra ${\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]$. Note that $^{\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]^{T}={\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}$. Therefore, we have an algebra homomorphism
(4.12) ${\mathbb{C}}[{\mathfrak{t}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}]_{\hat{0}}\rightarrow Z({\mathsf{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}).$
The functor $\pi_{*}:{\mathsf{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}})$ factors through a functor (still denoted by $\pi_{*}$) ${\mathsf{A}}\text{-}\mathrm{mod}\rightarrow\mathrm{Coh}({\mathfrak{g}}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$. By base change along ${\mathfrak{b}}_{S}\rightarrow{\mathfrak{g}}$, we obtain a functor
(4.13) ${\mathbb{V}}^{b}_{c}:\ {\mathsf{A}}_{S}^{b}\text{-}\mathrm{Mod}^{T}\rightarrow\mathrm{QCoh}^{T}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$
that is represented by the $({\mathbb{C}}[{\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}}],{\mathsf{A}}^{b}_{S})$-bimodule $\Gamma(\widetilde{{\mathfrak{g}}},{\mathcal{E}})\otimes_{{\mathbb{C}}[{\mathfrak{g}}]}{\mathbb{C}}[{\mathfrak{b}}_{S}]$.
On the other hand, by (2.8) there is an algebra homomorphism
@@
${\mathbb{V}}^{b}_{r}:\ {\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S}\rightarrow{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,0}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}\xrightarrow{{\mathsf{T}}^{-\rho}_{0}}{\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{Mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$
By the identification (3.8) and the compatibility of $\mathfrak{F}$$^{\prime}$ and ${\mathfrak{F}}^{b}$, the two maps (4.12) and (4.14) coincide under the equivalence ${\mathfrak{F}}^{b}$. Therefore, we obtain the following enhancement of (4.8)
(4.15) ${\mathsf{A}}_{S}^{b}\text{-}\mathrm{mod}^{T}$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,0}_{S}$$\mathrm{Coh}^{T}({\mathfrak{b}}_{S}\times_{{\mathfrak{t}}/W}{\mathfrak{t}})$${\mathfrak{U}}^{b}_{\zeta}\text{-}\mathrm{mod}^{\Lambda,-\rho}_{S\otimes_{{\mathbb{C}}[{\mathfrak{t}}/W]}{\mathbb{C}}[{\mathfrak{t}}]}.$$\scriptstyle{{\mathfrak{F}}^{b}}\scriptstyle{\simeq}$$\scriptstyle{{\mathbb{V}}^{b}_{c}}$$\scriptstyle{{\mathbb{V}}^{b}_{r}}$$\scriptstyle{{\mathfrak{G}}^{b}}\scriptstyle{\simeq}$